A jar contains 4 yellow marbles, 12 green marbles, and 11 black marbles. If one marble is selected at random, what is the probability that it is not yellow?
Answer by stanbon(48535)

A jar contains 4 yellow marbles, 12 green marbles, and 11 black marbles. If one marble is selected at random, what is the probability that it is not yellow?
----------------------
P(not yellow) = 1 - P(it is yellow)
= 1 - 4/27
= 23/27
=============
Cheers,
Stan H.